### v1

# colorlinks = true, linkcolor = blue, citecolor = blue, urlcolor = blue

Abstract
Object rearrangement is a significant task for collaborative robots, where they are directed to manipulate objects into a specified goal state. Determining the placement of objects is a major challenge that influences the efficiency of the rearrangement process. Most current methods heavily rely on pre-collected datasets to train the model for predicting the goal position and are restricted to specific instructions, which limits their broader applicability and effectiveness. In this paper, we propose a framework of language-conditioned object rearrangement based on the Large Language Model (LLM). Particularly, our approach mimics human reasoning by using past successful experiences as a reference to infer the desired goal position. Based on LLM’s strong natural language comprehension and inference ability, our method can generalise to handle various everyday objects and free-form language instructions in a zero-shot manner. Experimental results demonstrate that our methods can effectively execute the robotic rearrangement tasks, even those involving long sequential orders.
I INTRODUCTION
Robots have the potential to assist humans with various daily tasks, such as placing objects in desired locations, tidying up toys by gathering them into a designated container and performing household cleaning. These tasks can be categorised as object rearrangement, in which robots manipulate objects to achieve a specified goal state within a physical environment [1]. Effective object rearrangement involves several sub-tasks, including the recognition of object states within the environment, the inference of difference between the current and goal state, and the manipulation of objects accordingly. While there are many advanced solutions for object recognition and manipulation in computer vision, tactile sensing and robotic fields [2][3], we argue that there is a lack of intelligent methods that can perform human-like thinking to infer optimal object placement in rearrangement tasks. Specifically, achieving human’s ability to adapt to changing circumstances and guide flexible behaviour [4][5] by making use of similar past experiences on robots, still needs to be researched.
Currently, most existing methods attempt to learn a mapping from the current position to the goal position [6][7]. Firstly, a large dataset describing the goal position for the rearrangement needs to be collected. Then, the robot can be trained using the current position as the input to predict the goal position with an algorithm or a deep learning model. However, these methods rely heavily on the dataset, making them learn only specific patterns and human instructions from the dataset. When faced with a new environment containing different objects not included in the dataset or new instructions, these methods do not generalise effectively. Most recently, generative models like DALL-E, which is pre-trained on web-scale data, have been used to generate the goal poses in rearrangement to improve the generalisation ability [8]. However, these models can sometimes generate hallucinations, producing random objects and positions that require filtering in a further step.
Figure 1: Learn form the past. In our framework, the robot retrieves past experiences to find the most similar arrangement based on human instructions. By referencing previously successful arrangements, the robot can mimic human-like reasoning with LLMs, allowing it to infer the goal position for rearrangement more effectively.
When people rearrange items according to new instructions, it is natural for them to draw on their past successful experiences [9]. For example, “placing an apple on a plate” may remind them of a similar prior arrangement, such as “putting an orange in a bowl”. In this case, the experience of placing an orange in a bowl can serve as a template, providing scaffolding for the required robot trajectories and positions of both the orange and the bowl. This past experience then guides them in completing the new instructions. As a result, the efficiency of rearrangement tasks is significantly improved, as people distill and transfer knowledge from their previous experiences to new tasks.
Most recently, Large Language Models (LLMs) such as ChatGPT have significantly enhanced multiple fields within robotics, including code generation, task grounding, and human-robot interaction, thanks to their robust natural language understanding and zero/few-shot learning capabilities [10][11]. In this paper, we introduce a framework that links past experiences with new instructions by leveraging natural language comprehension (See Fig. 1). We then make use of an LLM to draw from similar past experiences and perform spatial reasoning to find the position for object placement in rearrangement tasks. As a result, our proposed method allows the robot to mimic human-like reasoning in object rearrangement, and our results demonstrate that the robots can execute rearrangement tasks efficiently, even in tasks with long sequential orders.
The contributions of this paper are summarised as follows:
  1. 1.
    We propose a novel framework based on LLMs that leverages past successful experiences to address the problem of language-conditioned object rearrangement.
  2. 2.
    Our method is not limited by the dataset and can be generalised to free-form natural language in a zero-shot manner.
  3. 3.
    By emulating human reasoning to learn from past experiences, our approach significantly enhances the performance of rearrangement tasks.
Figure 2: Illustration of the proposed framework. The robot uses SAM for visual perception and CLIP for semantic understandingidentify where and what objects are in the environment. The LLM then associate the most similar past experience with instruction and uses this similar experience as a reference. Finally, a prompt is created with spatial and semantic information, allowing the LLM to predict the goal position for rearrangement.
II Related works
In this section, we will first review works on predicting the goal positions in object rearrangement, followed by discussions of LLM’s applications in robotics. Finally, we will review the Visual Language Model (VLM) which connects visual perception and language instruction in robotics.
II-A Prediction of Goal Position in Rearrangement
The prediction of the goal position (That is the positions of each physical object in the environment) plays an important role in the object rearrangement task. The goal position can be represented in various ways such as through logical predicates, visual images or natural language instructions [1]. In [6][7], the spatial relations of the logical predicates, such as inside, on and next to are modelled between different objects to predict the goal position. Alternatively, for visual image representation, objects are manually arranged into goal states according to human preferences. In this situation, robots can then learn the latent space of user preferences from the image representation and manipulate the objects accordingly [12][13][14].
Additionally, some prior works treat position prediction as a searching task by selecting from a set of sampled poses using graphs. For example, in [15], a GNN is trained to understand the rules of the task, with its nodes representing entities relevant to the task such as objects and goals, which are used to select an appropriate position. In [16], an embodied agent is developed to detect out-of-place objects and infer plausible receptacles through a neural graph memory. However, the methods mentioned above typically require a large dataset that is manually collected for learning purposes, which can be both costly and time-consuming.
Compared to visual goal specification, language offers an intuitive way to describe the goal state. In [17][18], researchers detected objects within the environment and established spatial as well as semantic relationships between them. As a result, robots are able to relocate a specified object with respect to a designated anchor object towards fulfilling a set of given instructions. However, these methods concentrate on learning visual-semantic relationships and are unable to infer the goal position in a manner similar to human reasoning. In our framework, we leverage past successful experiences, similar to how humans do, and use the LLM to perform the spatial reasoning for the language-conditioned rearrangement, for the first of its kind.
II-B LLM and VLM in Language-conditioned Manipulation
Recently, LLMs have become an active area of research in the field of robotics due to their extensive internalised knowledge and chain of thought generation [11]. For example, in [19][16][20], language models are applied to decompose instructions into sequences of sub-steps. Based on the observed environment, the robot can execute the feasible sub-tasks to successfully achieve the instructions. Contrary to decomposing instructions, in [21], the LLM is applied to summarise the rules of arrangement from a small number of examples by leveraging its few-shot learning capabilities. In [10], the LLM is used to generate the planning strategy to enhance the stability of the system. In our framework, we employ the LLM to connect past experiences with new instructions and leverage it for spatial reasoning in rearranging objects, for the first time.
Most lately, Vision-Language Models (VLMs) have become a successful paradigm for aligning visual information with language, enabling robots to perceive the world in a multimodal manner. In [22], an end-to-end model based on VLMs is proposed. This model learns from large amounts of online data to create action commands that demonstrate robust generalisation capabilities in unseen environments. In [23], observations like images and state estimates are integrated into the language embedding space, facilitating more efficient inferences, particularly for sequential decision-making. However, these methods usually use large robotic data for training, which is a bottleneck because of limited data availability. To address this issue, [24] used pre-trained LLM and VLM to infer the affordances and constraints within the environment based on the language instructions provided. This approach enabled the execution of physical robot actions without relying on robotic data. In our framework, we apply the pretrained VLM to align the visual information with semantic information in a similar way and use the LLM together with past experience as a reference to perform spatial reasoning to estimate the goal position in a zero-shot manner.
III Methodology
The main objective of our framework is to leverage knowledge from previous arrangement experience and infer optimal positions for object placement conditioned on new instructions. The pipeline is shown in Fig. 2. First, the robots perform the vision-language grounding by using the semantic understanding provided by the Contrastive Language-Image Pre-Training Model (CLIP) and the vision precision from Segment Anything Model (SAM). This allows the robot to identify where and what the objects are in the environment. Then, we apply the LLM to associate the most relevant experience with the given instruction and use the spatial and semantic information from this similar experience as a reference. Next, we generate a prompt combining the spatial and semantic information of both the observed environment and the referenced experience. Based on this prompt, the LLM can effectively predict the position for object placement in rearrangement tasks, conditioned on the language instruction.
III-A Vision-language Grounding
For language-conditioned rearrangement, vision-language grounding, which is the aligning of each object with the corresponding semantic information, plays a fundamental role in the task. To predict the goal state of each object, we need to create an object-level representation, using the RGB observation of the environment from the camera. Specifically, we employ a pre-trained SAM for segmenting objects in RGB images. The model processes the image and point-based prompts as inputs and accordingly produces the object mask. By sampling single-point prompts across a grid on the image, SAM is able to generate a mask and bounding box for each point.
After generating the masks across a grid, semantic filtering is performed on the image to obtain the desired objects in the language instruction. In order to understand the objects to be manipulated, we first feed the LLM with language instruction and the prompt: “List the objects that are directly involved in the interaction described in the instruction”. In response, the LLM will identify and output the objects related to the instruction. For instance, if the instruction given is “put the apple next to the banana”, the LLM will output the objects involved as [“apple”, “banana”, “others”] where the “others” is given for the irrelevant objects. To align segmented objects with semantic information, the open-vocabulary image classifier CLIP is then used to match the masked images with corresponding names. Finally, an object representation is generated, which includes spatial information such as centre points, masks, bounding box, rotation, and the corresponding object names. Algorithm 3 demonstrates the core idea of the vision-language grounding with the observed RGB images.
III-B Spatial reasoning with retrieval augmentation generation
The key to object rearrangement is identifying the position where the target object should be placed. Due to the strong capabilities of language comprehension and internalised knowledge, LLM has great potential to become a general problem solver, extending beyond writing stories and essays. Based on the spatial information from observed images, such as the objects’ centre points, masks, bounding box, rotation, and semantic details, the LLM can be used to perform spatial reasoning to infer the objects’ target position for the rearrangement task.
Figure 3: Prompt engineering. The simplified prompt for LLM to perform spatial reasoning. It includes the spatial and semantic information from both the observed RGB image and a similar successful experience.
However, the dataset used to train the LLM is limited, as many LLMs are trained on corpora from the internet, books, and videos, but lack spatial information. Consequently, LLMs often struggle with spatial reasoning abilities and can not understand the shape accurately [25]. Inspired by the Retrieval Augmentation Generation (RAG) [26], we design a RAG system that allows the LLM to retrieve outer knowledge of past successful arrangement experiences and enhance its generated results. Specifically, this outer knowledge of successful rearrangement includes spatial and semantic information, along with the corresponding instructions. The system identifies the most similar arrangement from the outer knowledge as a reference. As shown in Fig. 2, the LLM scores the similarity between the new instruction and each instruction in past successful experiences. Detailed spatial and semantic information from the most closely matched experience is incorporated into the prompt to facilitate spatial reasoning. An example of a simplified prompt template is shown in Fig. 3. As a result, the LLM can predict the goal position more accurately during spatial reasoning.
{algorithm}
[…]
object_names
[…]
III-C Execute Pick and Place
Once the LLM identifies the object to be moved and the desired position for placement in the pixel coordinate, the robot can pick and place the object for the rearrangement. An RGB-D camera is mounted on the wrist of the robotic arm. Given the object’s initial and target locations in pixel coordinates, we can map these to their real-world coordinate using the depth information provided by the camera. We then employ inverse kinematics and a motion planner to find the path for the manipulation and move the object to complete the rearrangement task.
Figure 4: Examples of successful rearrangements. Four examples of successful rearrangements arranged by humans, with corresponding instructions.
Figure 5: Demonstration of results. The first row shows results from random object placement. The second displays objects arranged in a horizontal line with a certain gap. The third shows results from the proposed method without reference, and the fourth presents results using our proposed method.
IV Experiments and analysis
IV-A Robot Setup
For our experiment, we apply a UR5e robotic arm, which is equipped with an OnRobot RG2 gripper. An Intel RealSense D435 RGB-D camera is attached to the wrist of the UR5e robotic arm to obtain the RGB image and depth information. The robot is controlled via the Robot Operating System (ROS). In particular, we use the RRT-Connect algorithm for path planning.
IV-B Outer Knowledge of Successful Rearrangement
It is important to note that this outer knowledge of successful rearrangements is not used for training the model but rather serves as a reference to enhance the LLM’s spatial reasoning capabilities. The external knowledge from a successful experiment includes both spatial and semantic information, such as the name of each object, the object’s center point, its bounding box, rotation, and the corresponding instructions. 10 arrangements were created manually using representations of kitchen items. Specifically, the image of the successful human-arranged object rearrangement is captured from a top-down view. Based on the objects shown in the image, we use SAM to segment the corresponding masks using location prompts. From these masks, we derive the centre point, bounding box, and rotation of each object. Seven of the scenarios contain two objects, while the remaining three contain multiple objects. Some examples are shown in the Fig 4. Our framework also has the potential to expand external knowledge by recording the spatial information of objects during successful rearrangements performed by robots. However, in our experiments, we use only these 10 arrangements as outer knowledge to highlight how they improve the results.
IV-C User Study on Object Rearrangement
In our experiment, 15 arrangements are performed on 3 real-world scenes. The first scene consists of only an eggplant and a plate, with the robot manipulating a single object. The instructions for this scene are: (1) “put the eggplant on the right of the plate,” (2) “put the eggplant on the left of the plate,” (3) “put the eggplant in front of the plate,” (4) “put the eggplant behind the plate,” (5) “put the eggplant far away from the plate.” The second scene includes repetitive objects, i.e., two potatoes, to assess whether our method can effectively rearrange identical items and multiple objects. The instructions for the second scene are: (1) “put the potatoes on the plate,” (2) “put the potatoes beside the plate,” (3) “put one potato to the left of the plate and the other to the right,” (4) “put the potatoes far away from the plate,” (5) “put the potatoes together.”
In the third scene, our objective is to determine if our method can effectively rearrange objects in a long sequential order. This scene includes several objects, such as a carrot, a potato, a pineapple, an eggplant and a plate. The instructions for the third scene are: (1) “put the eggplant on the plate, then beside the plate,” (2) “put eggplant beside the plate, then beside the carrot,” (3) “put the eggplant beside the potato, then put the eggplant on the plate,” (4) “put the eggplant beside the carrot, then far away from the carrot,” (5) “put the eggplant on the right of the potato, then on the left of the pineapple.” In particular, we use the ChatGPT-4 [27] for spatial reasoning.
For the evaluation of object rearrangement, our objective is to measure if the final arrangement of objects matches the given instructions. Since our framework aims to mimic the human thinking method, a straightforward method is to have a user study to gather feedback from humans, following the approach in [16] [8]. Specifically, 10 participants were invited to evaluate whether the final real-world scene created by the robot meets the given instructions. Two of the participants are female, while the others are male, with ages ranging from 25 to 33. Participants rated the results on a scale from 1 to 10, where 1 indicates “not acceptable” and 10 indicates “highly acceptable”. All users were presented with the same set of images and the set-up for every method was similar. The participants evaluated the results from 8 different methods, providing a total of 1200 ratings.
As shown in Table I, we present a comparison of our results with other baseline methods. Since most existing approaches rely on large training datasets, they are not directly comparable to our method in a zero-shot manner. To address this, two heuristic-based baselines are selected for comparison. In the Random baseline, the robot picks objects and places them at random locations. In the Geometric baseline, the robot arranges the relevant objects in a horizontal line with a certain gap, ensuring their centre points are aligned at the same level. Moreover, we compare the results with an ablated structure that does not use the reference for spatial reasoning, which follows a similar structure as described in [21]. In [21], the LLM is used for categorisation, whereas we employ the LLM for spatial reasoning.
As shown in Table I, our proposed method achieves the highest average evaluation score of 9.09, outperforming other methods. This suggests that the robot can rearrange objects in a manner that closely resembles human-like behaviour. Furthermore, the results of our method without reference outperform those of heuristic-based baselines, highlighting the strong spatial reasoning capabilities of the LLM. Based on the LLM, the mean value increases by 1.16 while using the reference during spatial reasoning, indicating that our framework effectively transfers knowledge from prior successful experiences to facilitate more efficient object rearrangement. The examples in Fig. 5 also illustrate that the use of reference enhances the accuracy of object placement and enables a more reasonable gap between objects. However, the performance for single object rearrangement is lower than for other tasks. This is likely because the instructions include terms like “in front of” and “behind”, which make it more challenging for the LLM to accurately infer object positions. More results are available on our webpage at: https://sites.google.com/view/learnpast
IV-D Comparison between Different Language Models
In this experiment, we apply various language models to evaluate how their capabilities affect the prediction of goal positions in object rearrangement. Instead of using ChatGPT-4, we employ smaller LLMs to determine whether our proposed method can enhance performance even with models that require lower computational resources, thereby making their results comparable to those of more computationally intensive models.
Specifically, we use different backbone models Mistral-7B [28] and Llama3-8B [29] for spatial reasoning. These models are applied in both configurations, with and without reference, respectively. From Table II, we can observe that using a reference improves performance across all models in most tasks. In particular, the mistral model gains significant improvement from references in tasks involving multiple object rearrangement, with an increase of 3.64 in the score. However, Mistral and Llama3 show more variability with a higher standard deviation, which indicates that they are less consistent compared to ChatGPT-4. It can be seen ChatGPT-4 outperforms smaller models in almost every scenario, both with and without reference, but the performance gap narrows when reference is provided, compared to the other two models. Given that ChatGPT4 is significantly larger than Mistral-7B and Llama3-8B, the result suggests that the use of references can mitigate the limitations caused by decreased computational capacity.
TABLE I: Experimental results by using different methods
Methods Single object Multiple objects Sequential order Mean
Random 1.82±$\pm{1.61}$ 5.48±$\pm{2.50}$ 4.84±$\pm{3.40}$ 4.05
Geometric 2.98±$\pm{3.41}$ 4.76±$\pm{3.41}$ 4.80±$\pm{4.01}$ 4.18
w/o reference [21] 5.38±$\pm{3.74}$ 9.02±$\pm{1.82}$ 9.4±$\pm{1.31}$ 7.93
Ours 8.58±$\pm{\textbf{2.20}}$ 9.14±$\pm{\textbf{1.65}}$ 9.54±$\pm{\textbf{0.86}}$ 9.09
TABLE II: Experimental results by using different backbones
Model Single objects Multiple objects Sequential order Mean
mistral w/o reference 3.58±$\pm{3.39}$ 3.52±$\pm{3.01}$ 7.78±$\pm{2.90}$ 4.96
llama3 w/o reference 5.36±$\pm{3.36}$ 6.00±$\pm{4.21}$ 6.92±$\pm{3.77}$ 6.09
ChatGPT4 w/o reference 5.38±$\pm{3.74}$ 9.02±$\pm{1.82}$ 9.4±$\pm{1.31}$ 7.93
mistral w/ reference 5.02±$\pm{3.62}$ 7.16±$\pm{3.24}$ 9.14±$\pm{1.43}$ 7.11
llama3 w/ reference 7.74±$\pm{3.22}$ 6.74±$\pm{2.83}$ 8.56±$\pm{2.12}$ 7.68
chatGPT4 w/ reference 8.58±$\pm{\textbf{2.20}}$ 9.14±$\pm{\textbf{1.65}}$ 9.54±$\pm{\textbf{0.86}}$ 9.09
V Conclusion
In this paper, we propose an LLM-based framework for language-conditioned object rearrangement. By using past successful rearrangements as a reference, the robot can more accurately predict target positions for object placement, thereby mimicking the human reasoning process (see [4][5]). The results demonstrate that the use of references improve performance across multiple dimensions, including the rearrangement of a single object, multiple objects and tasks involving instructions with a sequential order. Additionally, the application of outer knowledge can help overcome the limitations caused by computational constraints, making it easier for practical applications in local deployment.
Nevertheless, the proposed method has some limitations that may affect its performance in real-world applications. Firstly, the scenario we designed operates primarily on a 2D surface, rather than in a 3D space, which limits its applicability in the physical world. Additionally, due to the spatial reasoning constraints of LLMs, we focus only on inferring the goal position but lack predicting the goal orientation for the arrangement. The objects used in the space are relatively sparse and this could pose a challenge in more clustered environments. Future improvements would involve extending the method to 3D rearrangement and addressing object rearrangement in clustered settings.